# Interpreting CLIP with Sparse Linear

### v1

@@ 7 Discussion
In this work, we show that the information contained in CLIP embeddings can be approximated by a linear combination of simple semantic concepts, allowing us to interpret representations with sparse, nonnegative dictionary learning. We propose SpLiCE , a method to transform the dense, uninterpretable embeddings of CLIP into human-interpretable sparse concept decompositions.
We empirically demonstrate that the improved interpretability of SpLiCE does not come at a cost to downstream performance, and provide three concrete use cases for SpLiCE : spurious correlation detection, model intervention and editing, and distribution shift monitoring, showcasing the benefits of using interpretable embeddings with known semantic content. We highlight that SpLiCE embeddings do not need to be used as a replacement for dense CLIP embeddings, but can also serve as a post-hoc interpretation of them.
@@
In this work, we only consider one word English concepts for our concept vocabulary; however, there exist many concepts that cannot be described under those constraints, such as “butter chicken”. In addition, we only consider linear decompositions, due to its convexity and potential added interpretability. Future work may consider using nonlinear decomposition solvers to achieve more accurate decompositions. SpLiCE also uses an $\ell_{1}$ penalty as the relaxation for $\ell_{0}$ regularization, but future work may consider alternative relaxations or even binary concept weights.
Broader Impact.
Similar to many works in the field of interpretability, our work provides greater understanding of the behavior of models, including but not limited to the broader implicit biases they perpetuate as well as mistakes made on individual samples. We believe this is particularly salient for CLIP, which is used in a variety of applications that are widely used. We hope that insights gained from such interpretability allow users to make more informed decisions regarding how they interact with and use CLIP, regardless of their familiarity with machine learning or domain expertise in the task they are using CLIP for. We also highlight that SpLiCE can be used as a visualization-like tool for exploring and summarizing datasets at scale, allowing for easier auditing of spurious correlations and biases in both datasets and models.
@@
[…]
Appendix A Relationship between cosine similarity and MSE optimization.
@@
Appendix B ADMM for batched on-device LASSO optimization.
As each decomposition requires solving a LASSO optimization problem, we implement the Alternating Direction Method of Multipliers (ADMM) algorithm in Pytorch over batches with GPU support for efficient decomposition of large scale datasets over large numbers of concepts Boyd et al. [2011]. In practice, ADMM achieves primal and dual tolerances of $41\mathrm{e}-$ 4 in fewer than $1000$ iterations on a batch size of $1024$
@@
Appendix C When do Sparse Decompositions Exist?
Theorem C.1.
Given Assumptions 4.1 - 4.4, CLIP image embeddings $f$ can be written as a sparse linear combination of text embeddings, i.e,
[…]
Proof.
$f(\mathbf{x}^{\text{img}})$ $=f\circ h^{\text{img}}(\omega,\epsilon)=f\circ h^{\text{img}}(\omega)=f\circ h^{\text{img}}\left(\sum_{i=1}^{k}\omega_{i}\mathbf{e}_{i}\right)=\sum_{i=1}^{k}\omega_{i}\underbrace{f\circ h^{\text{img}}(\mathbf{e}_{i})}_{\mathbf{c}^{\text{img}}_{i}}~{}~{}~{}~{}(\text{Assumption~{}4.2})$
Here we define $\mathbf{c}_{i}^{\text{img}}=f\circ h^{\text{img}}(\mathbf{e}_{i})$ as the ‘image’ concept basis vector; analogous to the text concept basis vector $\mathbf{c}_{i}^{\text{txt}}=g\circ h^{\text{txt}}(\mathbf{e}_{i})$ already defined. Thus Assumption 4.2 implies the existence of a sparse decomposition of $f$ in terms of ‘image’ concept vectors $\mathbf{c}^{\text{img}}_{i}$. Additionally, Assumption 4.1 ensures that this decomposition is sparse, as $\omega$ is sparse. So far, we have $f(\mathbf{x}^{\text{img}})=\mathbf{C}^{\text{img}}\omega~{}~{}~{}s.t.~{}~{}\|\omega\|_{0}\leq\alpha$.
From Assumption 4.3, the image concept vectors and text concept vectors are equal to each other, i.e, $\mathbf{c}_{i}^{\text{img}}=f\circ h^{\text{img}}(\mathbf{e}_{i})=g\circ h^{\text{txt}}(\mathbf{e}_{i})=\mathbf{c}_{i}^{\text{txt}}$. Finally, from Assumption 4.4, we have that the text concept vectors $\mathbf{c}_{i}^{\text{txt}}$ are given simply by word embeddings $g$ of individual words.
Appendix D Additional ImageNet Concept Histograms
We present concept histograms for the top seven concepts of five more ImageNet classes: {‘Face Powder’, ‘Feather Boa’, ‘Jack-O’-Lantern’, ‘Kimono’, ‘Dalmation’}, similar to Figure 6. These decompositions give insights both into the distribution of each class as well as some biases of CLIP. For example, the concepts “kitten” and “chihuahua” are present in the ‘Feather Boa’ decomposition, and were confirmed via manual inspection where we found 2 images of chihuahuas and 3 images of cats wearing boas in the dataset. For ‘Face Powder’, the concept “benefit” is the third most common concept, and it is indeed a common cosmetic brand name in the images. For the ‘Dalmation’ class, we see that the decompositions consists of concepts relating to dogs and black and white spots, which together make up the high-level concept of a dalmation. Finally, for the class ‘Kimono’, the concepts “doll”/“dolls” are the third and fourth most common, although all of the images in the ‘Kimono’ class were of real humans, not of dolls. This highlights an implicit bias in CLIP’s representations or in the descriptions of people wearing kimonos in CLIP’s training set.
Appendix E Additional Experiments
In this section we present additional experiments including (1) sanity checking the linearity of dense CLIP embeddings, (2) ablation studies on the mean centering operation, choice of concept vocabulary, and distribution of types of concepts highlighted by SpLiCE , (3) SpLiCE performance on probing tasks, (4) additional results for SpLiCE ’s performance on retrieval tasks, and (5) experiments on an alternative CLIP model from OpenAI with a ResNet50 visual backbone.
E.1 Sanity Checking CLIP’s Linearity
We begin by asking the following question to confirm the general linearity of CLIP embeddings: “if two inputs are concatenated, does their joint embedding equal the average of their two individual embeddings?”. For the image domain, we combine two images, $x_{a},x_{b}$, to form their composition $x_{ab}$ by placing $x_{a}$ in the top left quarter and $x_{b}$ in the bottom right quarter of a blank image. For the text domain, we simply append text $x_{b}$ to text $x_{a}$ to form $x_{ab}$. We then embed $x_{a},x_{b},x_{ab}$ with CLIP to get $z_{a},z_{b},z_{ab}$. Solving the equation $w_{a}*z_{a}+w_{b}*z_{b}=z_{ab}$ for scalar weights $w_{a},w_{b}$ then allows us to assess the linearity of $z_{a},z_{b},z_{ab}$. We report $w_{a},w_{b}$ and the cosine similarity between $\hat{z}_{ab}=[z_{a},z_{b}]\cdot[w_{a},w_{b}]$ and $z_{ab}$ in Table 2.
In general, we find that the composition of two inputs results in an embedding that is approximately equal to the average of the two input components, with $w_{a},w_{b}$ being very close to 0.5 across all datasets and for both modalities.
Table 2: Linearity of CLIP Embeddings.
$w_{a}$ $w_{b}$ cosine($\hat{z},z$)
ImageNet 0.48 $\pm$ 0.09 0.45 $\pm$ 0.09 0.76 $\pm$ 0.05
CIFAR100 0.45 $\pm$ 0.08 0.42 $\pm$ 0.08 0.75 $\pm$ 0.03
MIT States 0.48 $\pm$ 0.09 0.45 $\pm$ 0.09 0.76 $\pm$ 0.05
COCO Text 0.59 $\pm$ 0.12 0.47 $\pm$ 0.12 0.88 $\pm$ 0.04
E.2 Ablation Studies
Mean Centering.
In order to empirically check that the mean centering of images does not result in a loss of information, we decompose the img mean, $\mu_{img}$, that we used for all experiments. If we decompose it with uncentered concepts, the following concepts are highlighted: {“posed”, “added”, “pics”, “artist”, “caption”}. The decomposition with centered concepts results in the following concepts: {“closeup”, “posed”, “flickr”}. These concepts all seem to be generally related to images, with minimal other semantic information, suggesting that centering does not remove any discriminative semantic content of embeddings, but simply removes information about the modality.
Choice of Concept Vocabulary.
Concept Type Distribution.
In order to better understand any biases produced by the decomposition process or that CLIP itself has, we visualize the types of concepts most commonly activated across multiple datasets, labelling them by part of speech in Figure 11. We see that nouns are by far the most common concepts across datasets, indicating that both CLIP and the decompositions are highly object centric. Note that the low weight on verbs and adjective is due to far fewer concepts of those types being activated (low $l_{0}$ norm) as well as the weight upon those concepts being significantly smaller (low $l_{1}$ norm). We hypothesize that the information in many adjective and verbs can actually be encoded into the noun itself, resulting in this phenomenon. For example, the concept “lemon” is a more succinct form of “yellow” and “fruit”.
E.3 Performance of SpLiCE on Probing Tasks
@@
E.4 Additional Retrieval Experiments
We present additional results for the performance of SpLiCE embeddings on text-to-image and image-to-text retrieval tasks. We evaluate retrieval over various 1024 sample sub- sets of MSCOCO, and assess recall performance for the top-k closest embeddings of the opposite modality. We visualize $k=\{1,10\}$ in Figure 12 and $k=\{5\}$ in Figure 5. We find that our semantic concept dictionaries outperform all baselines when decomposition sparsity is high, but that dictionaries learned over images perform best for text to image retrieval when decompositions have greater than 30 nonzero concepts.
@@ E.4 Additional Retrieval Experiments
E.5 Experiments on Alternative CLIP Architecture

### v2

@@ 7 Discussion
In this work, we show that the information contained in CLIP embeddings can be approximated by a sparse, linear combination of simple semantic concepts, allowing us to interpret representations via sparse recovery. We propose SpLiCE , a method to transform the dense, uninterpretable embeddings of CLIP into human-interpretable sparse concept decompositions.
We empirically demonstrate that SpLiCE allows for an adjustable tradeoff on the interpretability-accuracy Pareto frontier, enabling users to decide the loss in performance they are willing to incur for interpretability. Furthermore, we find that the improved interpretability of SpLiCE allows for users to diagnose and fix model mistakes, ideally increasing the effectiveness and performance of the overall system using a VLM. We then provide concrete use cases for SpLiCE : spurious correlation detection and model intervention and editing, showcasing the benefits of using interpretable embeddings with known semantic content. We highlight that SpLiCE embeddings can serve as post-hoc interpretations of CLIP embeddings and can even replace them to ensure full transparency.
Broader Impact.
Similar to many works in the field of interpretability, our work provides greater understanding of the behavior of models, including but not limited to the broader implicit biases they perpetuate as well as mistakes made on individual samples. We believe this is particularly salient for CLIP, which is used in a variety of applications that are widely used in practice at this moment. We hope that insights gained from such interpretability allow users to make more informed decisions regarding how they interact with and use CLIP, regardless of their familiarity with machine learning or domain expertise in the task they are using CLIP for. We also highlight that SpLiCE can be used as a visualization-like tool for exploring and summarizing datasets at scale, allowing for easier auditing of spurious correlations and biases in both datasets and models.
@@
In this work, we use a large, overcomplete dictionary of one- and two-word concepts, however future work may wish to expand this dictionary or learn a dictionary over tokens (in discrete language space), to capture concepts with more than two words. This may also reduce the size of the dictionary and improve computation time. We note that this dictionary was constructed by looking at token frequency in the LAION-5B dataset, which has its own biases and may not correctly capture all the salient concepts that CLIP encodes. Despite this, we find that SpLiCE  performs well on a variety of tasks while outperforming state-of-the-art concept dictionaries (Fig. 5, Appendix Fig. 13) and thus we believe LAION is a good dataset to generate a concept vocabulary from. We also note that this vocabulary can be easily modified by practitioners to consider additional concepts as needed for specific use cases. Finally, SpLiCE also uses an $\ell_{1}$ penalty as the relaxation for $\ell_{0}$ regularization, but future work may consider alternative relaxations or even binary concept weights.
@@
This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, FAI-2040880, and research awards from Google, JP Morgan, Amazon, Adobe, Harvard Data Science Initiative, and the Digital, Data, and Design (D3) Institute at Harvard. AO is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2140743, and UB is funded by the Kempner Institute Graduate Research Fellowship. The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.
Appendix
Summary of Appendix Results
  • A. Further Details on the Method
    • A.1. When do Sparse Decompositions Exist?
    • A.2. Relationship between cosine similarity and MSE optimization
    • A.3. ADMM for batched on-device LASSO optimization
    • A.4. Effect of Modality Alignment
    • A.5. Experimental Details
  • B. Additional Results
    • B.1. User Study for Human Interpretability
    • B.2. Performance of SpLiCE on Probing Tasks
    • B.3. Performance of SpLiCE on Retrieval Tasks
    • B.4. Additional Zero-Shot Results
    • B.5. Additional ImageNet Concept Histograms
    • B.6. Additional Case Study: Detecting Spurious Correlations
    • B.7. Additional Case Study: Spurious Correlation Intervention
    • B.8. Additional Case Study: Distribution Shift Monitoring
    • B.9. Additional Case Study: Distribution Shift Monitoring
    • B.10. Checking the Interpretability of Negative Concepts
    • B.11. Understanding the Image Mean for Modality Alignment
    • B.12. Choice of Concept Vocabulary
    • B.13. Concept Type Distribution
    • B.14. Experiments on Alternative CLIP Architecture
Appendix A Further Details on the Method
A.1 When do Sparse Decompositions Exist?
Proposition 1.
Given Assumptions 1-5, CLIP image embeddings $f$ can be written as a sparse linear combination of text embeddings, i.e,
$f(\mathbf{x}^{\text{img}})=\mathbf{C}^{\text{txt}}\mathbf{w};~{}~{}s.t.~{}~{}\|\mathbf{w}\|_{0}\leq\alpha$
where $\mathbf{w}\in\mathbb{R}_{+}^{k}$, and $\mathbf{C}^{\text{txt}}\in\mathbb{R}^{d\times k}$, which is the text concept dictionary defined previously.
Proof.
$f(\mathbf{x}^{\text{img}})=$ $f\circ h^{\text{img}}(\omega,\epsilon)=f\circ h^{\text{img}}(\omega)~{}~{}~{}~{}(\text{Assumption~{}2})$
$=$ $f\circ h^{\text{img}}\left(\sum_{i=1}^{k}\omega_{i}\mathbf{e}_{i}\right)=\sum_{i=1}^{k}\omega_{i}\underbrace{f\circ h^{\text{img}}(\mathbf{e}_{i})}_{\mathbf{c}^{\text{img}}_{i}}~{}~{}~{}~{}(\text{Assumption~{}3})$
Here we define $\mathbf{c}_{i}^{\text{img}}=f\circ h^{\text{img}}(\mathbf{e}_{i})$ as the ‘image’ concept basis vector; analogous to the text concept basis vector $\mathbf{c}_{i}^{\text{txt}}=g\circ h^{\text{txt}}(\mathbf{e}_{i})$ already defined. Thus Assumption 2 implies the existence of a sparse decomposition of $f$ in terms of ‘image’ concept vectors $\mathbf{c}^{\text{img}}_{i}$. Additionally, Assumption 1 ensures that this decomposition is sparse, as $\omega$ is sparse. So far, we have $f(\mathbf{x}^{\text{img}})=\mathbf{C}^{\text{img}}\omega~{}~{}~{}s.t.~{}~{}\|\omega\|_{0}\leq\alpha$.
From Assumption 4, the image concept vectors and text concept vectors are equal to each other, i.e, $\mathbf{c}_{i}^{\text{img}}=f\circ h^{\text{img}}(\mathbf{e}_{i})=g\circ h^{\text{txt}}(\mathbf{e}_{i})=\mathbf{c}_{i}^{\text{txt}}$. Finally, from Assumption 5, we have that the text concept vectors $\mathbf{c}_{i}^{\text{txt}}$ are given simply by word embeddings $g$ of individual words.
A.2 Relationship between cosine similarity and MSE optimization.
@@ A.1 When do Sparse Decompositions Exist?
A.3 ADMM for batched on-device LASSO optimization.
As each decomposition requires solving a LASSO optimization problem, we implement the Alternating Direction Method of Multipliers (ADMM) algorithm in Pytorch over batches with GPU support for efficient decomposition of large scale datasets over large numbers of concepts [64]. In practice, ADMM achieves primal and dual tolerances of $41\mathrm{e}-$ 4 in fewer than $1000$ iterations on a batch size of $1024$ We present an empirical comparison beterrn LASSO and ADMM in 6, where we find both methods to be approximately equivalent.
@@
Figure 6: Comparison of ADMM (maroon) and LASSO (yellow) for solving the SpLiCE objective on zero shot accuracy (left) and cosine reconstruction (right) on CIFAR100. Both methods are approximately equal.
A.4 Effect of Modality Alignment
We take MSCOCO images and captions, embed them with CLIP, and compare the cosine similarity between modalities and inter-modality. Before mean-centering and renormalizing, the similarity within modalities is high, with an average of around 0.3. This indicates that the image and text embeddings do not span the entire unit-sphere but rather lie on two cones. However, the similarity across modalities has an average concentrating around zero, indicating that these two cones are non-overlapping. However, after mean-centering and normalizing, we observe that the average cosine similarity for images, text, and between images and text becomes zero and the modalities are aligned.
A.5 Experimental Details
All experiments are able to be performed on a single A100 GPU to run fast inference with CLIP. After embedding the concept dictionary, all computation can be performed on a CPU. Code is made available at https://github.com/AI4LIFE-GROUP/SpLiCE.
Appendix B Additional Results
B.1 User Study for Human Interpretability
We present results from a user study in 8 to assess the human interpretability of SpLiCE. We base our study off of that performed by [23] to evaluate Label-Free Concept Bottleneck Models (LF-CBMs). We benchmark our method against LF-CBMs and IP-OMP [22]. We provided users with twenty randomly chosen, correctly predicted images from ImageNet and explanations from two different methods comprising the top six most important concepts for every image. We then asked users to evaluate and compare the different concept-based explanations for (1) their relevance to the provided image inputs, (2) their relevance to model predictions, and (3) their informativeness on Likert scales from 1 to 5. We found that users significantly preferred explanations generated by SpLiCE to the two baselines for relevance to the images and informativeness, with significance determined via a one-sample two-sided t-test and a threshold of p=0.01. We also highlight that our method is able to produce similar/better concept decompositions, in terms of human interpretability, than the baselines without needing to train a classification probe or use class labels for concept mining, both of which are computationally expensive. This user study was ruled exempt by our institution’s IRB, as no risks were posed to the users. Participants were able to opt out at any time, and no questions were asked regarding the participants themselves.
Figure 8: Results of a user study evaluating SpLiCE, LF-CBM, and IP-OMP in the style of the user study from LF-CBM. Overall, we find that explanations generated by SpLiCE are deemed more relevant to the image, relevant to the prediction, and more informative than prior methods.
B.2 Performance of SpLiCE on Probing Tasks
@@
B.3 SpLiCE Performance on Retrieval Tasks
We test the performance of SpLiCE embeddings on text-to-image and image-to-text retrieval tasks. We evaluate retrieval over various 1024 sample subsets of MSCOCO, and assess recall performance for the top-k closest embeddings of the opposite modality for $k=\{1,5,10\}$. We find that our semantic concept dictionaries outperform all baselines when decomposition sparsity is high, but that dictionaries learned over images perform slightly better for text to image retrieval when decompositions have greater than 30 nonzero concepts.
@@ B.3 SpLiCE Performance on Retrieval Tasks
B.4 Additional Zero-Shot Results
We present additional results comparing SpLiCE reconstructed vectors and CLIP embeddings on the Caltech101 [65], SUN397 [66], STL10 [67], and VOC2007 [68] datasets in 5. We use SpLiCE decompositions with sparsities of 20-35, and we find that they are comparable to the unaltered CLIP embeddings.
Table 5: Additional zero-shot accuracy on baselines from the CLIP paper, for decompositions of sparsity 20-35. Note that at human-interpretable levels of sparsity, we see a minor drop in performance.
Caltech101 SUN397 STL10 VOC 2007
CLIP Reported 0.88 0.63 0.97 0.83
CLIP Implemented 0.90 0.67 0.96 0.92
SpLiCE 0.86 0.66 0.96 0.83
We further explore the performance of SpLiCE decompositions in the limit as they approach the sparsity of the baseline CLIP embeddings (512). We find that SpLiCE completely recovers CLIP zero-shot accuracy at this limit, as shown in 6.
Table 6: Zero shot performance at sparsity 512. Note that SpLiCE completely recovers baseline CLIP zero shot accuracy.
CIFAR100 MITStates Imagenet
CLIP Baseline 0.750 0.469 0.552
SpLiCE (512) 0.768 0.474 0.552
B.5 Additional ImageNet Concept Histograms
We present concept histograms for the top seven concepts of five more ImageNet classes: {‘Face Powder’, ‘Feather Boa’, ‘Jack-O’-Lantern’, ‘Kimono’, ‘Dalmation’}, similar to Figure 10. These decompositions give insights both into the distribution of each class as well as some biases of CLIP. For example, for the class ‘Face Powder’, the concept “benefit" is the fifth most common concept, and it is indeed a common cosmetic brand name in the images. For the ‘Dalmation’ class, we see that the decompositions consists of concepts relating to dogs and black and white spots, which together make up the high-level concept of a dalmation. Finally, for the class ‘Kimono’, the concept “doll" is the seventh most common, although all of the images in the ‘Kimono’ class were of real humans, not of dolls. This highlights an implicit bias in CLIP’s representations or in the descriptions of people wearing kimonos in CLIP’s training set.
B.6 Additional Case Study: Detecting Spurious Correlations
We present an additional case study for detecting spurious correlations in CIFAR100. In particular, we look at the prevalence of the spurious concept “desert” in the classes ‘camel’ and ‘kangaroo’ in Figure 11. We observe that camels are more frequently pictured in the desert, creating a spurious signal that may be leveraged by downstream classifiers. This figure provides an additional example of how we can understand biases and trends in data with SpLiCE  decompositions.
Figure 11: Distribution of “Desert" concept in ‘Camel’ and ‘Kangaroo’ classes of CIFAR100.
B.7 Additional Case Study: Spurious Correlation Intervention
We further test the ability of SpLiCE to enable intervention on intermediate representations and linear classifiers by attempting to remove information pertaining to spurious signals. In particular, we consider the Waterbirds dataset [63], which spuriously correlates landbirds with land backgrounds, resulting in trained classifiers performing poorly on waterbirds on land. We thus remove information about whether or not birds are on land backgrounds by ablating concept weights on “bamboo”, “forest”, “hiking”, and “rainforest” as well as any bigrams containing the word “forest,” as shown in Table 7. This significantly improves worst-case subgroup performance for waterbirds on land from 0.48 to 0.60.
For both this experiment and the intervention on CelebA described in the main paper, we train linear probes using the LogisticRegressionClassifier module in scikit-learn using an $\ell_{1}$ penalty.
Table 7: Evaluation of intervention on spurious correlations for Waterbirds dataset. Removing information about land backgrounds improves worst-case subgroup performance.
Landbirds
on land
Waterbirds
on land
Linear Probe 0.98 0.48
Intervention
Probe
0.97 0.60
B.8 Additional Case Study: Distribution Shift Monitoring
We present a final case study using SpLiCE to monitor distribution shift. This can help identify differences between training and inference distributions or evaluate how a continually sampled dataset changes over time. In this experiment we consider the Stanford Cars dataset [69], which contains photos of cars from 1991 to 2012, including their make and year labels. By decomposing photos of cars from each year, we can view how the distribution changed yearly. We visualize the weights of the concepts “convertible" and “yellow" from our decompositions, as well as the actual percentage of cars from each year that were convertibles or yellow in Figure 12. Note the right-hand y-axis, corresponding to the weight of the given concept $c_{i}$ over the sum of the weights of all concepts $\sum_{i}c_{i}$, does not have a meaningful unit of measure or scale. We find that the trends in the groundtruth concept prevalence generally closely match that of the predicted/decomposed concepts, allowing us to visualize which years convertibles or yellow cars were popular or out-of-distribution with respect to other years. Most notably, we see that SpLiCE picks up on the out-of-distribution rise in popularity of brightly colored sports cars in the early 2000s.
B.9 Additional Case Study: Distribution Shift Monitoring
To further verify that SpLiCE allows for identification and tracking of distribution shift, we study the Waterbirds dataset, which is known to have differently balanced train, valodation, and test splits. To identify distribution shifts, we can simply look at the norm of the difference between the class decompositions of the two classes for each splot, as shown in 8. We find that the validation and test splits are much more similar than the training and validation splits or the training and test splits, which can be verified by the construction process of the Waterbirds dataset.
Table 8: Study of the differences in distributions between train, validation, and test splits of Waterbirds. The validation and test splits are much more similar to each other than they are to the train split.
Train, Val Train, Test Val, Test
Class Landbird 0.0182 0.0182 0.005
Class Waterbird 0.0229 .0188 0.009
We also find that the most weighted concept in the ‘landbird’ class of the train split is “bamboo” but the corresponding weight for “bamboo” in the ‘waterbird’ class is much lower. The “bamboo” concept weight for both classes and all splits is shown below, where we see that the validation and test splits are very similar and mostly evenly balanced, whereas the train split is highly unbalanced.
Table 9: Study of the prevalence of the concept “bamboo” in the different classes and splits of Waterbirds.
Train Val Test
Class Landbird 0.0196 0.010 0.010
Class Waterbird 0.0007 0.008 0.008
B.10 Checking the Interpretability of Negative Concepts
We take a set of 71 concept-antonym pairs from the MIT States dataset and embed the terms in CLIP. With and without concept centering, we observe that these concept-antonym pairs have an average cosine similarity well above -1, indicating that CLIP does not place antonyms in opposite directions, as shown in 10. Next, we take our concept dictionary and prepend “not" to all of the words and compare the average cosine similarity between concept and not-concept pairs. Similarly, we observe that with and without centering, concept and not-concept pairs are highly similar. Note that the average similarity for true pairs of images and text in MSCOCO is less than the similarity between concepts and not-concepts with and without centering.
Table 10: Evaluation of the similarity of antonyms and negative concepts in CLIP.
Pairwise Cosine Similarity
(without concept centering)
Pairwise Cosine Similarity
(with concept centering)
Concept and antonym 0.7176 $\pm$ 0.1109 0.1366 $\pm$ 0.2197
Concept and “not” concept 0.8661 $\pm$ 0.0498 0.6130 $\pm$ 0.0498
B.11 Understanding the Image Mean for Modality Alignment
In order to empirically check that the mean centering of images does not result in a loss of information, we decompose the img mean, $\mu_{img}$, that we used for all experiments. If we decompose it with uncentered concepts, the following concepts are highlighted: {“closeup", “flickr", “posed"}. The decomposition with centered concepts results in the following concepts: {“flickr", “posed", “pics", “angle view", “last post"}. These concepts all seem to be generally related to images, with minimal other semantic information, suggesting that centering does not remove any discriminative semantic content of embeddings, but simply removes information about the modality.
B.12 Choice of Concept Vocabulary
B.13 Concept Type Distribution
In order to better understand any biases produced by the decomposition process or that CLIP itself has, we visualize the types of concepts most commonly activated across multiple datasets, labelling them by part of speech in Figure 14. We see that nouns are by far the most common concepts across datasets, indicating that both CLIP and the decompositions are highly object centric. Note that the low weight on verbs and adjective is due to far fewer concepts of those types being activated (low $l_{0}$ norm) as well as the weight upon those concepts being significantly smaller (low $l_{1}$ norm). We hypothesize that the information in many adjective and verbs can actually be encoded into the noun itself, resulting in this phenomenon. For example, the concept “lemon" is a more succinct form of “yellow" and “fruit".
B.14 Experiments on Alternative CLIP Architecture